# Reasoning-Aware Proxy Reward Model using Process Mining

### v2

@@ Abstract
Recent advances in sparse reward policy gradient methods have enabled effective reinforcement learning (RL) for language models post-training. However, for reasoning tasks such as mathematical problem solving, binarized outcome rewards provide limited feedback on intermediate reasoning steps. While some studies have attempted to address this issue by estimating overall reasoning quality, it remains unclear whether these rewards are reliable proxies for the quality of stepwise reasoning. In this study, we consider reasoning as a structured process and propose TACReward, the reward model that can be seamlessly integrated into sparse reward policy gradient methods without additional human annotation costs or architectural modifications. TACReward aggregates stepwise structural deviations between teacher and policy reasoning using process mining techniques, producing a scalar output reward range of $[0,1]$ to indicate reasoning quality. Experiments on multiple mathematical reasoning benchmarks demonstrate that integrating the TACReward into sparse reward frameworks encourages the policy model to improve the structural quality of reasoning. Consequently, this leads to consistent performance improvements over existing sparse reward frameworks. Our code and model are publicly available at GitHub and HuggingFace
Large Reasoning Model, Reward, Policy Optimization, Process Mining
@@ 1 Introduction
Recent advances in post-training methods based on RL (Ouyang et al., 2022) have established a new paradigm for enhancing the alignment and reasoning capabilities of Large Reasoning Models (LRMs) (Zhang et al., 2025). Traditional policy gradient methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and its variants, have achieved significant success. To improve efficiency, Reinforced Leave-One-Out (RLOO) (Ahmadian et al., 2024) introduced a leave-one-out baseline to reduce the variance without a dedicated critic. GRPO (Shao et al., 2024) and GSPO (Zheng et al., 2025) have shifted this paradigm by adopting sparse reward signals and removing the critic network and Generalized Advantage Estimation (GAE).
@@
Despite these advances, policy gradient methods that rely on sparse reward signals remain challenging to apply to reasoning tasks, such as mathematical problem solving. This difficulty arises from their dependence on verifiable outcome rewards (i.e., reward sparsity)  (Cui et al., 2025). This reliance on the binary and rule-based signals often produces uniform reward signals when responses within a problem group exhibit similar correctness. Consequently, this yields minimal differentiation and weakened learning gradients (Zhang and Zuo, 2025). One potential solution involves providing step-level labels such as Process Reward Models (PRMs) (Lightman et al., 2023). However, adopting such dense reward within a sparse reward framework is challenging without architectural modifications, potentially diluting the advantage of stability (Sullivan, 2025). Furthermore, defining step-level labels is inherently challenging, and human annotation remains costly (Jaech et al., 2024; Lightman et al., 2023).
@@
Recent studies have proposed sparse reward frameworks that consider reasoning quality without annotation (Cui et al., 2025; Fan et al., 2025; Yang et al., 2025; Zhan et al., 2025), most of which are based on GRPO. However, these methods often rely on an indirect estimation or approximation of the overall reasoning process quality. This makes it challenging to consider the resulting reward signal as indicator of the quality of the individual reasoning steps. Consequently, concerns remain regarding the reliability of resulting reward signals, as it remains ambiguous whether improvements in cumulative rewards genuinely reflect enhanced logical maturity or merely the over-optimization of the surrogate metrics. Therefore, while ensuring that step-level logical integrity is crucial for authentic reasoning, an inherent dilemma arises from its sparse reward nature.
@@
To address these challenges, this study proposes a Trace, Alignment, and Check Reward (TACReward), a novel reasoning-aware reward model that can be seamlessly integrated into sparse reward policy gradient methods for LRM post-training. TACReward treats the reasoning of LRMs as a structured process rather than a monolithic output. The core idea is that TACReward encourages the policy model to improve its degree of alignment with the teacher model reasoning process, which is considered more mature than the policy model. The degree of alignment was computed as the sparse final reward signal in the range 0 and 1 by aggregating stepwise structural deviations between the teacher and policy models. Therefore, TACReward provides proxy signals for the quality of the reasoning process from a structural alignment perspective. To this end, process mining (Van Der Aalst, 2016), a set of techniques specialized in analyzing data recording the execution of processes (i.e., event logs), is adapted and extended. Notably, these procedures require no additional human annotation.
Our proposed TACReward was evaluated using multiple mathematical benchmarks (Hendrycks et al., 2021; He et al., 2024; MAA, n.d.; Lewkowycz et al., 2022; KICE, 2025; Liu et al., 2025b). The experimental results showed that integrating TACReward into sparse reward policy gradient methods can enhance performance over RLOO, GRPO, and GSPO across all benchmarks. Notably, GSPO + TACReward achieved an average relative accuracy improvement of 89.2 %.
@@ 2.1 Event Log & Process Mining
Figure 1: Example of mathematical reasoning event log with a single case id and the corresponding process model (i.e., petri-net) inferred from it. The boxes in the map represent activities, and the arrows indicate the process flow. The circles denote the places, which are used to model the state of the process.
An event log (Van Der Aalst et al., 2011) is hierarchically structured data that records the execution of a process. The formal definitions of events, traces, and logs are as follows:
Definition 2.1 (Event, Trace, and Log).
E is the event universe. An event is a tuple $e=\left(c,a,t,d_{1},\dots,d_{m}\right)\in E$, where $c$ is the case id; $a$ is the activity; $t$ is the timestamp, and $d_{1},\dots,d_{m}$ are the additional attributes. A trace $\sigma=\left\langle e_{1},e_{2},\dots,e_{n}\right\rangle\in E^{*}$ is a finite, non-empty sequence of events such that a trace contains all and only the events in the same case id. An event log $\mathbb{L}$ is a multiset of traces.
Process mining (Van Der Aalst, 2016) is a set of techniques that uses event logs to discover process models (Leemans et al., 2013), check the conformance between event log and process model (Carmona et al., 2022), and enhance or extend the models. Figure 1 presents an example of an event log and the process model derived from it. Various discovery algorithms have been proposed ranging, from foundational Alpha Miner (Van Der Aalst et al., 2011) to more robust approaches, such as inductive miners (Leemans et al., 2013).
@@ 2.3 Sparse Reward Policy Gradient Methods for LRMs
The policy gradient methods (Sutton et al., 1999) with sparse rewards maximizes expected reward for the responses generated by the policy LRM $\pi_{\theta}$:
@@
where $R(x,y)$ is the reward function that provides a scalar reward signal based on the quality of the response $y$ to query $x$. Using the score function estimator (Williams, 1992), the policy gradient can be expressed as
@@
It is often convenient to define the (Monte Carlo) advantage estimator (Sutton et al., 1998) as: $\hat{A}(x,y)\triangleq R(x,y)-b(x)$, such that Equation 2 can be expressed as $\hat{A}(x,y)$.
@@ 2.4 Group-based Policy Gradient Methods for LRMs
Recent sparse-reward post-training methods often sample multiple responses for the same query $x$ (Shao et al., 2024; Zheng et al., 2025). For each $x$, we draw a group of $G$ responses
@@
3 TAC: Trace, Align, and Check
The proposed TACReward treats the reasoning of an LRM as a structured process rather than a monolithic output. By leveraging and Extending Techniques From process mining, TACReward evaluates the logical integrity of a reasoning trace without requiring step-level human annotations or architectural changes to the policy optimization framework. TACReward encourages a policy model to improve the degree of alignment with the teacher model’s reasoning process, which is logically more mature than the policy model. DeepSeek R1 (Guo et al., 2025) was used as the teacher model because of its strong reasoning capabilities. The core concept of TACReward is illustrated in Figure 2.
In process mining literature, conformance checking aims to evaluate how a process model can explain an entire event log that contains multiple traces (Van Der Aalst, 2016). In the proposed TACReward, this paradigm was extended to operate on individual traces. TACReward can be categorized into three main steps: 1) Trace: Formalizing the reasoning trace of both the policy and teacher models. 2) Align: Examines the alignment between the policy and reference traces. 3) Check: Evaluate how well the reasoning process of the policy model conforms to that of the teacher model.
@@ 10 Apply Theorem 20 End Problem
The computation of the TACReward begins by prompting the policy model to generate a reasoning process for a given problem $x$. For each problem, the model produces $G$ candidate responses, $\{y_{i}\}_{i=1}^{G}$, each containing a sequence of reasoning steps. The steps are enclosed within <think><think> tags (see Section B.1 for the generation prompt). Subsequently, these responses were formalized into reasoning traces by segmenting them into discrete steps and formatting them to conform to an event- log schema. Each fundamental units of reasoning represents a distinct activity. The taxonomy of the 20 activities is adopted based on related studies and mathematical problem-decomposition frameworks (Qin et al., 2025; Berti et al., 2025; Polya, 2014; Ritter et al., 2019). A complete list of the activities is provided in Table 1.
To extract the traces, we employed DeepSeek-V3.2 (Liu et al., 2025a), which is a powerful general-purpose language model. Let $\pi_{\psi}$ denote a general purpose model: The policy traces are formalized as:
@@
where $\sigma_{i}$ is the trace corresponding to the $i$-th response $y_{i}$. Similarly, the reference trace $\pi_{\phi}$ is formalized as:
@@
where the teacher model generates a single response for each problem. The prompt used to formalize the trace is provided in Section B.2.
@@ 3.2 Align: Examine the Alignment
In this step, each policy trace is transformed into a process model using the Inductive Miner (IM) algorithm (Leemans et al., 2013). Given the policy traces $\{\sigma_{i}\}_{i=1}^{G}$ and reference trace $\sigma^{ref}$, we discover a process model for each policy trace by treating $\{\sigma_{i}\}$ as single-trace event log:
@@
where $\mathcal{M}_{i}$ denotes the discovered process model (e.g., a Petri net) corresponding to the $i$th policy trace.
To establish an alignment (Van der Aalst et al., 2012) between a policy process model $\mathcal{M}_{i}$ and the reference trace $\sigma^{ref}$, we relate the moves in the model to the moves in the trace. Let $\Sigma$ be the set of activity labels and let $\gg$ denote a ”no-move” (i.e., misalignment) symbols. Given model $\mathcal{M}_{i}$, let $\beta(\mathcal{M}_{i})\subseteq\Sigma^{*}$ be the set of all possible complete executions sequences in $\mathcal{M}_{i}$.
Definition 3.1 (Alignment).
@@
To illustrate this, consider the reference trace $\sigma^{ref}=\langle A,D,B,E,G\rangle$ and the process model in Figure 1 as the policy process model $\mathcal{M}_{i}$. Two example alignments are as follows:
[…]
where $\gg$ indicates a misalignment (a log-only move or model-only move), where each column represents a vertical movement. For example, In $\gamma_{1}$, the first move $(A,A)$ is a synchronous move. The the second move $(D,\gg)$ indicates that the reference trace executes $D$ whereas the model does not. In $\gamma_{2}$, the second move $(\gg,C)$ indicates that: that the model executes $C$ that is absent from the reference trace.
To quantify the discrepancy between the reference traces $\sigma^{ref}$ and the behavior allowed by the policy process model $\mathcal{M}_{i}$, a finite cost is assigned to each move in an alignment.
Definition 3.2 (Cost Function).
Let $(a_{ref},a_{m})$ be a move in alignment $\gamma\in\Gamma(\sigma^{ref},\mathcal{M}_{i})$. The move cost function $\delta_{m}$ is defined as
@@
where $w_{L}(\cdot)\geq 0$ and $w_{M}(\cdot)\geq 0$ are non-negative weights. We use $w_{L}(\cdot)=w_{M}(\cdot)=1$ unless stated otherwise. These three cases correspond to the following: Synchronous move, move on log, and move on model, respectively.
Subsequently, the total cost of the alignment, $\delta(\gamma)$, is computed by summing the move costs over all the moves in the alignment of $i$-th model:
@@ 3.3 Check: Evaluate the Conformance
Through the alignment step, an optimal alignment $\gamma_{i}^{*}$ is established between each policy process model $\mathcal{M}_{i}$ and the reference trace $\sigma^{ref}$. Based on this alignment, we quantify the conformance score $r^{TACReward}_{i}$, which serves as the TACReward output for each policy response. We computed the conformance by using the F1-score of fitness and precision (Van der Aalst et al., 2012).
Fitness. Fitness measures how well the process model explains the observed behavior in the the reference traces (Van Der Aalst, 2016). To obtain stable normalization, the finite the worst-case deviation cost is defined as:
$\delta_{worst}(\sigma^{ref},\mathcal{M}_{i})$ (11)
$=\sum_{e\in\sigma^{ref}}w_{L}(e)$ $\quad+\min_{\sigma^{\beta}\in\beta(\mathcal{M}_{i})}\sum_{e^{\prime}\in\sigma^{\beta}}w_{M}(e^{\prime})$
the cost of $\sigma^{ref}$ with only log moves, plus the minimum cost to complete the model using only the model moves. Subsequently, the fitness is computed as
@@
Precision. The precision assesses how often actions that are not present in the trace occur in a process model (Van Der Aalst, 2016). Let $en_{M}(e)$ be the set of enabled activities in model $\mathcal{M}_{i}$ in the state immediately before event $e$ occurs, and let $en_{L}(e)$ be the set of activities observed in $\sigma^{ref}$ in the same context (i.e., the same prefix immediately before $e$). The precision is computed as:
@@
Output of TACReward. Based on the computed fitness and precision, the output reward $r^{TAC}_{i}$ is defined as
@@
where $r^{TAC}_{i}\in[0,1]$ denotes the degree of conformance between the reasoning process of the policy model and in the teacher model for the $i$th response. TACReward can be seamlessly integrated into sparse-reward policy gradient methods by directly adding its output as the original task reward. Detailed integration procedures are provided in Appendix C.
3.4 TAC as a Proxy for Step-level Reward
The alignment $\gamma_{i}^{*}$ decomposes the discrepancy between the policy and teacher reasoning traces into individual moves, each incurring local cost $\delta_{m}(\cdot)$. The total deviation cost performs implicit credit assignment at the reasoning transition level, even though the final reward is scalar. A log-only move $(a_{ref},\gg)$ indicates that the policy model omits the reasoning activity that is present in the teacher’s trace and corresponds to a missing logical step. Conversely, the model-only move $(\gg,a_{m})$ reflects the introduction of redundant or irrelevant reasoning activities that are not supported by the teacher’s process.

### v3

# Rewarding Structural Conformance of Reasoning using Process Mining

@@ Abstract
Recent advances in sparse reward policy gradient methods have enabled effective reinforcement learning (RL)-based language model post-training. However, for reasoning tasks such as mathematical problem solving, binarized outcome rewards provide limited feedback on intermediate reasoning steps. While some studies have attempted to address this issue by estimating overall reasoning quality, it remains unclear whether these rewards are reliable proxies for the quality of stepwise reasoning. In this study, we consider reasoning as a structured process and propose TACReward, the reward model that can be seamlessly integrated into sparse reward policy gradient methods without additional human annotation costs or architectural modifications. TACReward aggregates stepwise structural deviations between teacher and policy reasoning using process mining techniques, producing a scalar output reward range of $[0,1]$ to indicate reasoning quality. Experiments on multiple mathematical reasoning benchmarks demonstrate that integrating the TACReward into sparse reward frameworks encourages the policy model to improve the structural quality of reasoning. Consequently, this leads to consistent performance improvements over existing sparse reward frameworks. Our code and checkpoints are publicly available at GitHub and HuggingFace.
@@ 1 Introduction
Recent advances in post-training methods based on RL [25] have established a new paradigm for enhancing the alignment and reasoning capabilities of Large Reasoning Models (LRMs) [47]. Traditional policy gradient methods, such as Proximal Policy Optimization (PPO) [30] and its variants, have achieved significant success. However, PPO incurs substantial computational and memory overhead from maintaining a separate critic network and often suffers from training instability when scaled to large reasoning models. To improve efficiency, Reinforced Leave-One-Out (RLOO) [1] introduced a leave-one-out baseline to reduce the variance without a dedicated critic. GRPO [31] and GSPO [48] have shifted this paradigm by adopting sparse reward signals and removing the critic network and the Generalized Advantage Estimation (GAE).
@@
Despite these advances, sparse reward policy gradient methods remain challenging to apply to reasoning tasks, such as mathematical problem solving due to their dependence on verifiable outcome rewards (i.e., reward sparsity) [6]. The reliance on binary and rule-based signals often produces uniform reward signals when responses within a problem group exhibit similar correctness. Consequently, this yields minimal differentiation and weakened learning gradients [46]. One potential solution involves providing step-level labels such as Process Reward Models (PRMs) [17]. However, adopting such dense reward within a sparse reward framework is challenging without architectural modifications, potentially diluting the advantage of stability [33]. Furthermore, defining step-level labels is inherently challenging, and human annotation remains costly [17].
@@
Recent studies have proposed sparse reward frameworks that consider reasoning quality without annotation [6, 8, 43, 45], most of which are based on GRPO. However, these methods often rely on an indirect estimation or approximation of the overall reasoning process quality. This makes it challenging to consider the resulting reward signal as indicator of the quality of the individual reasoning steps. Consequently, it remains unclear whether improvements in cumulative rewards reflect genuine gains in logical maturity or mere over-optimization of surrogate metrics. Therefore, while ensuring that step-level logical integrity is crucial for authentic reasoning, an inherent dilemma arises from sparse reward nature.
We start from the observation that solutions to a reasoning problem can differ substantially at the surface level yet share a common structural skeleton of cognitive activities. In mathematical problem solving, for instance, one solution may proceed by substitution and another by direct computation, but both typically follow a sequence such as Formulate Strategy $\rightarrow$ Apply Formula $\rightarrow$ Simplify Expression $\rightarrow$ Verify. This observation suggests that reasoning quality can be characterized structurally, in terms of the sequence and composition of such activities, rather than the specific content of individual reasoning steps.
Figure 1: TACReward separates correct from incorrect reasoning. Distributions of TACReward, computed against the reasoning of DeepSeek-R1 (671B) [9] as the reference, for correct (green) and incorrect (purple) solutions on DeepMath-103k [12]. Four models ranging from 1.5B to 70B parameters are evaluated. Correct responses consistently receive a higher degree of alignment than incorrect ones, and the separation is statistically significant for all models (U-test [23], $p<10^{-10}$).
@@
Building on this view, we propose Trace, Alignment, and Check Reward (TACReward), a novel reasoning-aware reward model that can be seamlessly integrated into sparse reward policy gradient methods for language model post-training (i.e., the final reward signal is in the range of 0 to 1) with no additional human annotation cost. In TACReward, the reasoning of LRMs is treated as a structured activity sequence (i.e., process) rather than a monolithic output and the policy model is encouraged to improve its degree of alignment with the teacher model’s reasoning, which is considered more mature than the policy model. To this end, process mining [38], a set of techniques specialized in analyzing data recording the execution of processes, is adapted and extended. Our preliminary investigation shows that TACReward, computed against the reference, assigns systematically higher scores to responses with correct final answers than to those with incorrect ones, with a statistically significant separation between the two distributions (Figure 1).
@@ 2.1 Event Log & Process Mining
Figure 2: Example of mathematical reasoning event log with a single case id and the corresponding process model inferred from it. The boxes in the map represent activities, and the arrows indicate the flow. The circles denote the places, which are used to model the state of the process.
An event log [36] is hierarchically structured data that records the execution of a process. The formal definitions of events, traces, and logs are as follows:
Definition (Event, Trace, and Log).
E is the event universe. An event is a tuple $e=\left(c,a,t,d_{1},\dots,d_{m}\right)\in E$, where $c$ is the case id, $a$ is the activity, $t$ is the timestamp, and $d_{1},\dots,d_{m}$ are the additional attributes. A trace $\sigma=\left\langle e_{1},e_{2},\dots,e_{n}\right\rangle\in E^{*}$ is a finite, non-empty sequence of events such that a trace contains all and only the events in the same case id. An event log $\mathbb{L}$ is a multiset of traces.
Process mining [38] is a set of techniques that uses event logs to discover process models, check the conformance between event log and process model [3], and enhance or extend the models. Figure˜2 presents an example of an event log and the process model derived from it. Various discovery algorithms have been proposed ranging from foundational Alpha Miner [36] to more robust approaches, such as inductive miners [15].
@@ 2.3 Sparse Reward Policy Gradient Methods for LRMs
The policy gradient methods [35] with sparse rewards maximizes expected reward for the responses generated by the policy LRM $\pi_{\theta}$:
@@
where $R(x,y)$ is the reward function that provides a scalar reward signal based on the quality of the response $y$ to query $x$. Using the score function estimator [41], the policy gradient can be expressed as
@@
It is often convenient to define the (Monte Carlo) advantage estimator [34] as: $\hat{A}(x,y)\triangleq R(x,y)-b(x)$, such that Equation˜2 can be expressed as $\hat{A}(x,y)$.
@@ 2.4 Group-based Policy Gradient Methods for LRMs
Recent sparse-reward post-training methods often sample multiple responses for the same query $x$ [31, 48]. For each $x$, we draw a group of $G$ responses
@@
3 Method
The proposed TACReward treats the reasoning of an LRM as a structured process rather than a monolithic output and aims to encourage a policy model to improve the degree of alignment with the teacher model’s reasoning process. To this end, techniques of process mining are leveraged and extended. In process mining literature, conformance checking aims to evaluate how a process model can explain an entire event log that contains multiple traces [38]. We extend this paradigm to operate on individual traces. Specifically, TACReward can be categorized into three main steps: 1) Trace: Formalizing the reasoning trace of both the policy and teacher models. 2) Align: Examines the alignment between the policy and reference traces. 3) Check: Evaluate how well the reasoning process of the policy model conforms to that of the teacher model.
The proposed TACReward is illustrated in Figure˜3.
@@ 10 Apply Theorem 20 End Problem
The computation of the TACReward begins by prompting the policy model to generate a reasoning process for a given problem $x$. For each problem, the model produces $G$ candidate responses, $\{y_{i}\}_{i=1}^{G}$, each containing a sequence of reasoning steps (see Section˜D.1 for the generation prompt). Subsequently, these responses were formalized into reasoning traces by segmenting them into discrete steps and formatting them to conform to an event log schema. Each structural units of reasoning represent a distinct activity. The taxonomy of the 20 activities is adopted from related studies and mathematical problem-decomposition frameworks [27, 2, 26, 29]. A complete list of the activities is provided in Table˜1. Table˜8 provides the ablation study on the taxonomy.
To extract the traces, DeepSeek-V3.2 [18], a general-purpose language model, is employed. Let $\pi_{\psi}$ denote a general-purpose model, the policy traces are formalized as:
@@
where $\sigma_{i}$ is the trace corresponding to the $i$-th response. Similarly, the reference trace is:
@@
where the teacher model generates a single response for each problem. The prompt used to formalize the trace is provided in Section˜D.2.
@@ 3.2 Align: Examine the Alignment
In this step, each policy trace is transformed into a process model using the Inductive Miner (IM) algorithm [15]. Given the policy traces $\{\sigma_{i}\}_{i=1}^{G}$ and reference trace $\sigma^{ref}$, we discover a process model for each policy trace by treating $\{\sigma_{i}\}$ as a single-trace event log:
@@
where $\mathcal{M}_{i}$ denotes the discovered process model corresponding to the $i$th policy trace.
To establish an alignment [37] between a policy process model $\mathcal{M}_{i}$ and the reference trace $\sigma^{ref}$, we relate the moves in the model to the moves in the trace. Let $\Sigma$ be the set of activity labels and let $\gg$ denote a "no-move" (i.e., misalignment) symbol. Given model $\mathcal{M}_{i}$, let $\beta(\mathcal{M}_{i})\subseteq\Sigma^{*}$ be the set of all possible complete executions sequences in $\mathcal{M}_{i}$.
Definition (Alignment).
@@
To illustrate this, consider the reference trace $\sigma^{ref}=\langle A,D,B,E,G\rangle$, and the process model in Figure˜2 as the policy process model $\mathcal{M}_{i}$. For clarity, two example alignments $\gamma_{1}$ and $\gamma_{2}$ are provided in Figure˜4. In an alignment, $\gg$ indicates a misalignment (a log-only move or model-only move), where each column represents a vertical movement. When the same activity is executed in both the reference trace and the model, it is called a synchronous move.
$\begin{aligned} \gamma_{1}&=\left|\begin{array}[]{c|c|c|c|c|c}A&D&B&\gg&E&G\\ \hline\cr A&\gg&B&D&E&G\end{array}\right|\\[8.0pt] \gamma_{2}&=\left|\begin{array}[]{c|c|c|c|c|c}A&\gg&D&B&E&G\\ \hline\cr A&C&D&\gg&E&G\end{array}\right|\end{aligned}$
Figure 4: Two example alignments $\gamma_{1}$ and $\gamma_{2}$. In $\gamma_{1}$, the first move $(A,A)$ is a synchronous move. The second move $(D,\gg)$ indicates that the reference trace executes $D$ whereas the model does not. In $\gamma_{2}$, the second move $(\gg,C)$ indicates that the model executes $C$ that is absent from the reference trace.
To quantify the discrepancy between the reference traces and the behavior allowed by the policy process model $\mathcal{M}_{i}$, a finite cost is assigned to each move.
Definition (Cost Function).
Let $(a_{ref},a_{m})$ be a move in alignment $\gamma\in\Gamma(\sigma^{ref},\mathcal{M}_{i})$. The move cost function $\delta_{m}$ is defined as:
@@
where $w_{L}(\cdot)$ and $w_{M}(\cdot)$ are non-negative weights and $w_{L}(\cdot)=w_{M}(\cdot)=1$ unless stated otherwise.
Finally, the optimal alignment $\gamma_{i}^{*}$ for the $i$-th model is obtained by minimizing the total cost $\delta(\gamma)$, which is the sum of the costs over all moves in the alignment:
@@
Appendix˜C provides the computational complexity analysis of finding $\gamma_{i}^{*}$.
@@ 3.3 Check: Evaluate the Conformance
Through the alignment step, an optimal alignment $\gamma_{i}^{*}$ is established between each policy process model $\mathcal{M}_{i}$ and the reference trace $\sigma^{ref}$. Based on this alignment, we quantify the conformance score $r^{TAC}_{i}$, which serves as the TACReward output for each policy response. We compute the conformance by using the F1-score of fitness and precision [37].
Fitness. Fitness measures how well the process model explains the observed behavior in the reference traces [38]. To obtain stable normalization, the finite worst-case deviation cost is defined as:
$\delta_{worst}(\sigma^{ref},\mathcal{M}_{i})=\sum_{e\in\sigma^{ref}}w_{L}(e)$ $\quad+\min_{\sigma^{\beta}\in\beta(\mathcal{M}_{i})}\sum_{e^{\prime}\in\sigma^{\beta}}w_{M}(e^{\prime})$ (11)
where the cost of $\sigma^{ref}$ with only log moves, plus the minimum cost to complete the model using only the model moves. Subsequently, the fitness is computed as
@@
Precision. The precision assesses how often actions that are not present in the trace occur in a process model [38]. Let $en_{M}(e)$ be the set of enabled activities in model $\mathcal{M}_{i}$ in the state immediately before event $e$ occurs, and let $en_{L}(e)$ be the set of activities observed in $\sigma^{ref}$ in the same context (i.e., the same prefix immediately before $e$). The precision is computed as:
@@
Output of TACReward. Based on the computed fitness and precision, the output reward $r^{TAC}_{i}$ is:
@@
where $r^{TAC}_{i}\in[0,1]$ denotes the degree of conformance between the reasoning process of the policy model and that of the teacher model for the $i$th response. TACReward can be seamlessly integrated into sparse-reward policy gradient methods by directly adding its output as the original task reward. Detailed integration procedures are provided in Appendix˜A.